Radiant heat makes it impossible to stand close to a hot lava flow. Calculate the rate of heat loss by radiation from 1.00 m2 of 1110°C fresh lava into 36.2°C surroundings, assuming lava's emissivity is 1.
Answer #1


rate of heat loss= sEA(T- Ts) Ans The A = 1m2 , Ts= E Emissivity 36-20 + 273= 304-2 K TR=lava temperature = l10°+273 13 83 K